QUESTION THREE
a. Calculate the rate of capital gain or loss on a ten-year zero-coupon bond for which the
Interest rate has increased from 10% to 20%. The bond has a face value of $1,000.
{10 Marks)
b. Why have financial intermediaries become so important in the economy? Do you think
they will survive in their present form? {10 Marks
1
Expert's answer
2021-09-20T18:57:08-0400

a)Price of Zero-coupon bond today = "\frac{maturity value}{(1+r)^n}"


"= \frac{1000}{(1+0.1)^{10}}" "=385.84"


"= \frac{1000}{(1+0.2)^{9}}" "= 193.81"


Rate of capital "\frac{gain}{loss}" = "\frac{P1-P0}{P0}"


"= \frac{193.81-385.84}{385.84}" = "-49.73"

= -49.73%


b) Financial intermediaries help in widening of the financial services to customers. It is usually hard for other people to access main financial institutions, which through intermediaries, they are easily facilitated.

The intermediaries will still survive in their present forms since there are some areas where people are unable to access services.